# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v2

@@ III.2.1 Quantum experiment details
This section provides technical details about the experiments carried out on the quantum devices. The circuits that we use come from the truncation of two layers of the LUCJ circuits (as described in Sec. III.1). For the quantum runs, the circuit parameters are not optimized and instead are obtained from a classical CCSD calculation, as described in Sec. III.1. Note that the circuits themselves are still classically challenging to simulate.
@@
Table 2 lists the different molecular species, together with the number of qubits used to encode the $|\Psi\rangle$ that generates samples of electronic configurations. The same table also provides details about the total circuit depths, and numbers of one- and two-qubit gates, as well as the IBM device that was used to run the experiments. The specific qubit layouts are also provided in Table 2 and depicted in Fig. S9.
@@
Table 2 shows the 95% confidence interval for that fraction. Recall that the measurement outcomes that correspond to configurations with the correct particle number are used in the setup phase of the configuration recovery procedure. Table 2 also shows the fraction of sampled configurations with the correct particle number if samples were collected from the uniform distribution over configurations of length $M$,
@@
As shown in the Table, the two probabilities are statistically distinguishable for all the experiments we carried out. Our experiments used twirled readout error mitigation (ROEM) [102] to mitigate errors arising from qubit measurement, and dynamical decoupling (DD) [103, 104, 105, 106] to mitigate errors arising from quantum gates. We employed the implementation of ROEM and DD available on the Runtime library of Qiskit [107], through the Sampler primitive. DD is implemented by sequences of $X$ control pulses, whose effect is to protect qubits from decoherence due to low-frequency system-environment coupling. Here, we applied sequences of two $X$ pulses (as in Ramsey echo experiments) to idle qubits.
@@
Table 2: Details of hardware simulations. For each active space (column 1 from the left), we list the details of the LUCJ quantum circuit executed on hardware, specifically: its number of qubits (q) both for the Jordan-Wigner encoding (JW) and the total number of qubits used including auxiliary ones (Tot.) (column 2) and count of quantum operations (circuit depth $d$, number of $\mathsf{CNOT}$ and single-qubit $\mathsf{u}$ gates, column 3), the device used (column 4) and the qubit layout chosen (column 5, where $[a\rightarrow b]$ indicates the shortest path between qubits $a$ and $b$ in the device topology, also visible in Fig. S9). In columns 6 and 7 we report a 95% confidence interval for the fraction $p_{N}^{\mathrm{hw}}$ of configurations with the correct particle number and the corresponding value for configurations with uniform probability distribution (see the text for the definition of these quantities). In column 8, we list the number of noisy configurations sampled from quantum hardware.
Figure S9: Left to right: schematics of the devices used to carry out experiments for the N2 molecule with 6-31G basis, the N2 molecule with cc-pVDZ basis, the [2Fe-2S] cluster, and the [4Fe-4S] cluster, using local unitary cluster Jastrow (LUCJ) quantum circuits. Qubits used in the calculation are shown in red (for qubits associated with $\alpha$ spin-orbitals), blue for qubits associated with $\beta$ spin-orbitals), and green (for auxiliary qubits).
III.2.2 Classical resource details
The SCI-based eigenstate solver used by HPC quantum estimator is the one implemented in PySCF [26, 63] for the N2 (6-31G, cc-pVDZ), and [2Fe-2S] systems. For the [4Fe-4S] molecule we use the SCI-based eigenstate solver implemented in DICE [38, 35] because of its ability to run across multiple computer nodes.
@@ III.2.2 Classical resource details
IV Quantum circuits that can sample more efficiently than classical Selected Configuration heuristics
In this section, we show that there are LUCJ circuits that can sample electronic configurations more efficiently than some classes of classical heuristics, yielding equal or better energies than classical heuristics at fixed subspace dimensions.
For our study we consider a particular flavor of SCI, Heat bath Configuration Interaction (HCI) [34], and the [2Fe-2S] cluster, representative of strongly correlated species (20 orbitals $\rightarrow$ 40+ qubits). The circuit under consideration is the same LUCJ circuit that was used for the hardware experiments shown in the main text, with the same Heavy-Hex connectivity.
We first run HCI with different values of the HCI selection cutoff $\epsilon_{1}$ [81], as shown in Fig. S10. We observe that the larger and more restrictive values of $\epsilon_{1}$ do not allow HCI to reach the DMRG energy reference. Instead, it converges to the first excited state (S1) energy. Decreasing the value of $\epsilon_{1}$ allows the subspace dimension to become larger and improve the energy beyond that first excited state, at the cost of a significant increase in subspace dimension, which results in a higher computational cost. We note that for $\epsilon_{1}>8\cdot 10^{-5}$ it takes two or more HCI iterations to leave the first excited state. From the largest HCI wave function (last point for $\epsilon_{1}=5\cdot 10^{-5}$), we repeatedly remove the electronic configurations whose wave function amplitude is below a given threshold and compute the energy in the resulting subspace, providing a collection of energy-dimension pairs that we label as “upper bound to optimal” in Fig. S10. The removal of tens of millions of configurations (resulting in significantly smaller subspace dimensions) does not significantly deteriorate the quality of the ground-state approximation. Eventually, when the subspace dimension becomes smaller than $50\cdot 10^{6}$, the deterioration of the energy becomes more notable. These results indicates that HCI does not perform an optimal search of the relevant electronic configurations. Consequently, much smaller subspaces exist that yield comparable energy values. In this particular molecule, HCI needs to explore and perform diagonalizations in subspaces much larger than the optimal search.
After assessing that there exist subspaces of smaller dimension that yield equivalent energy estimations to HCI, we optimize an LUCJ circuit to produce samples in the identified subspaces with high probability. The optimal parameters are found in a two step optimization workflow. First, we optimize the Kullback–Leibler divergence between samples drawn from the LUCJ circuit and the amplitudes of 58 M-dimensional ground-state wave function estimation (green arrow in Fig. S10). Then, the circuit parameters are further fine-tuned to minimize the energy using a differential-evolution strategy.
With the samples produced by the optimized LUCJ circuit, we proceed to evaluate the energy of resulting subspaces of varying dimensionality, including the 58 M-dimensional subspace that the circuit was optimized for. We observe good agreement between the LUCJ subspaces and those that are an upper bound to the optimal ones. It is worth noting that despite being optimized for a 58 M-dimensional subspace, the resulting circuit produces subspaces of outstanding quality of dimensionalities both much smaller and larger. This result highlights the high ability of the LUCJ circuit to explore the space of electronic configurations.
Figure S10: Comparison between HCI and optimized LUCJ circuits in the efficiency of exploration of relevant electronic configurations to describe the ground state of [2Fe-2S]. The panels show the energy estimation as a function of subspace dimension of the diagonalization required to reach the corresponding energy. The dots connected by arrows show the trajectory of HCI in the energy-dimension plane. Each dot is labelled by the iteration it corresponds to. Different panels correspond to HCI calculations carried out with different values of $\epsilon_{1}$, as indicated in the legend. Brown triangles show energies and subspace dimensions obtained after truncating the most accurate HCI wave function obtained by repeatedly removing the electronic configurations of lowest amplitude. The green triangles show the energies for different subspace dimensions defined by samples obtained from an optimized LUCJ circuit. The green arrows show the subspace dimension that the circuit was optimized for. The horizontal solid and dashed lines indicate the DMRG estimates for the S0 and S1 eigenstates.
The active space considered for the [2Fe-2S] cluster is small enough for HCI to be able to find the relevant bitstrings by exploring and performing diagonalizations in subspaces significantly larger than the optimal ones. This shows that in some problem instances, the diagonalization itself is not the runtime bottleneck. Instead, the bottleneck is a sub-optimal proposal of electronic configurations. Larger and more strongly correlated systems will pose more challenges for the classical heuristics. It is these systems where quantum circuits can play a role.
V Experimental results
@@
This section contains the results from the experiments that are run on the quantum devices, as described in the previous section. We show a collection of investigations that are complementary to those shown in the main text.
@@
We first visualize the prediction individual wave function components obtained by the HPC quantum estimator with configuration recovery in a system size where we have access to the exact ground state wave function amplitudes (N2 in the 6-31G basis).
@@
V.1 Visualization of the wave function amplitudes
@@
Panel (b) in Fig. S11 shows a comparison between the exact ground state wavefunction amplitudes and the wavefunction amplitudes obtained by the estimator, for different points in the dissociation curve. The agreement is exceptional for the larger wavefunction amplitudes and slowly deteriorates in the description of the tails of the wavefunction. We also observe that for larger bond lengths, where higher levels of static correlation are present, the agreement in the tails is worse than for bond lengths closer to the equilibrium geometry. This is to be expected for two reasons. First, for larger bond lengths the wavefunction amplitudes are less concentrated than close to equilibrium. As discussed in Sec. II.2, the accuracy of the estimator using a SCI-based eigenstate solver deteriorates as the wavefunction becomes less concentrated. The second reason is that for bond lengths close to equilibrium, the wavefunction has a strong mean field character, but not for larger bond lengths. This directly impacts the shape of the reference spin-orbital occupancy ${\bf{n}}$ that is used to recover configurations. When the wavefunction is mean-field dominated, ${\bf{n}}$ takes the shape of a sharp step function shape with all spin-orbital average occupancies close to 0 or 1. One possible effect of static correlations is to soften the step function, moving average occupancies of some spin-orbitals away from 0 or 1. The recovery of configurations used in this work is more effective when the components of ${\bf{n}}$ are close to 0 or 1.
@@
In summary, the HPC quantum estimator with a SCI-based eigenstate solver and configuration recovery produces accurate representations of the ground state wavefunction (on experimental data), even at an individual amplitude level. Less concentrated wavefunctions tend to challenge the procedure more, both because the accuracy of the SCI solver is decreased, and because the configuration recovery becomes less effective due to softer profiles of ${\bf{n}}$.
@@
Figure S11: Visualization of the wavefunctions produced the the HPC quantum estimator estimator in the dissociation of N2 (6-31G), and the comparison against the exact wavefunctions. (a) Shows the energy and energy error as a function of the bond length for the estimator with configuration recovery run from measurement outcomes from the quantum device. Classical approximate methods are shown for reference. (b) Shows the wavefunction amplitudes $|c_{{\bf{x}}}|^{2}$ for all possible electronic configurations ${\bf{x}}$. The configurations in the horizontal axis are ordered according to the magnitude of the amplitudes in the exact wavefunction. The blue solid lines correspond to the exact ground state amplitudes while the orange dots correspond to the wavefunction amplitudes obtained from the estimator with configuration recovery. Different points in the dissociation curve are shown and labelled A, B, C, D, following the same encoding as in panel (a).
V.2 Scaling and runtime
@@
We begin this section by studying the improvement of the accuracy of the HPC quantum estimator with the number $d$ of determinants used for the subspace expansion and diagonalization. It is clear that larger $d$ yields more expressivity in the variational wavefunction, since we are increasing the possible size of the support of the state. It is also clear that increasing $d$ increases the HPC quantum estimator runtime since both the projection and diagonalization processes become more computationally intensive. To show the improvement of the accuracy with $d$ we consider the dissociation of N2 (cc-pVDZ), where the measurement outcomes are generated by the quantum device. The estimator is applied to $10^{5}$ measurement outcomes and we take $K=10$ batches of samples and report $\min_{k}\left(E^{(k)}\right)$ for the potential energy surface for different values of $d$. Ten self-consistent iterations are considered for the recovery of configurations.
@@
Fig. S12 shows the potential energy surfaces obtained by the estimator with different values of $d$, the number of determinants used for the subspace expansion and diagonalization. Increasing the value of $d$ improves the results both qualitatively and quantitatively. For the smaller values of $d$, the potential energy surface shows unphysical oscillations when the bond length is large. For larger values of $d$, not only the energy values are lower, but also the oscillations in the energy as a function of the bond length are significantly decreased in amplitude. Improvement of the ground state approximation with an increasing amount of computational resources is also observed in the energy-variance analyses of the main text (Fig. 4), and the supplementary materials (Figs. S21 and S22).
@@
Figure S12: Improvement of the estimator accuracy (with configuration recovery) by increasing the number of determinants for the subspace expansion and diagonalization. Potential energy surface of the N2 molecule (cc-pVDZ) obtained from the HPC quantum estimator estimator using measurements outcomes from the quantum device. The energies reported correspond to the lowest energy amongst the K batches of configurations, i.e. ${\min_{k}\left(E^{(k)}\right)}$. Different subspace dimensions $d$ are considered, as indicated in the legend. Different classical methods are shown for reference.
@@
To conclude this section, we study the dependence of the runtime and accuracy of the estimator in the [4Fe-4S] system as a function of the amount of classical compute. The classical compute can be increased by increasing the size of the subspace dimension $d$, or by collecting more batches of configurations $K$. The DICE library allows to distribute the subspace diagonalization across a number of independent computer nodes. It is expected that more classical nodes will yield faster runtimes per diagonalization. For this experiment the Heron device Montecarlo is used, with $\left|\mathcal{X}\right|=3163742$ collected measurement outcomes, and $p_{N}^{hw}=5.4\cdot 10^{-4}$.
@@
Figure S14: Accuracy and runtime of the HPC quantum estimator for the [4Fe-4S] iron-sulfur cluster. (a) Runtime of a single subspace projection and diagonalization as a function of the number of nodes that are used to distribute the calculation. Different colors correspond to different subspace dimensions. The smallest number of nodes per curve corresponds to the limit where the calculation runs out of memory on a single node. The subspace projection and diagonalization are handled by the DICE library. The horizontal line shows a fixed runtime of 1.44 hours for the different subspace sizes. (b) Statistical convergence of the HPC quantum estimator as the number of batches of samples $K$ is increased. The vertical axis shows the the average minimum energy obtained by a fixed number of batches of samples. Each batch of samples is run on fixed number of nodes and the diagonalization for the different batches are run in parallel. The number of nodes for the diagonalization of a single batch is chosen to obtain a fixed runtime of 1.44 hours for all of the different subspace dimensions $d$. The horizontal axis shows the total number of nodes corresponding to different numbers of batches. The largest number of batches corresponds to $K=100$.
V.3 Effect of orbital optimizations in the dissociation of N2 (cc-pVDZ)
We apply orbital optimizations (OO) to improve the accuracy of the HPC quantum estimator, in the study of the dissociation of N2 (cc-pVDZ). The estimator with configuration recovery and its orbital-optimized counterpart are applied to the same set of noisy measurement outcomes $\tilde{\mathcal{X}}$ obtained from the quantum device. We use $K=10$ batches of samples in each case and $N_{\textrm{SCF}}=10$ iterations of the alternation between the optimization of $\kappa$ and running of the eigenstate solver to update $c_{\bf{x}}^{(k)}$. After each $c_{\bf{x}}^{(k)}$ update, $5000$ iterations of gradient descent with momentum are used to optimize $\kappa$. Before applying orbital optimizations, the configuration recovery procedure is run for 10 iterations on the reference basis (molecular orbitals). Various sizes $d$ of the batches are considered.

### v3

@@ 3.1 Quantum and classical computational resources
This section provides technical details about the experiments carried out on the quantum processors. The circuits that we use come from the truncation of two layers of the LUCJ circuits (as described in Sec. Quantum circuits). For the quantum runs, the circuit parameters are not optimized and instead are obtained from a classical CCSD calculation, as described in Sec. Quantum circuits. Note that the circuits themselves are still classically challenging to simulate.
@@
Table S1 lists the different molecular species, together with the number of qubits used to encode the $|\Psi\rangle$ that generates samples of electronic configurations. The same table also provides details about the total circuit depths, and numbers of one- and two-qubit gates, as well as the IBM quantum processor that was used to run the experiments. The specific qubit layouts are also provided in Table S1 and depicted in Fig. S8.
@@
Table S1 shows the 95% confidence interval for that fraction. Recall that the measurement outcomes that correspond to configurations with the correct particle number are used in the setup phase of the configuration recovery procedure. Table S1 also shows the fraction of sampled configurations with the correct particle number if samples were collected from the uniform distribution over configurations of length $M$,
@@
As shown in the Table, the two probabilities are statistically distinguishable for all the experiments we carried out. Our experiments used twirled readout error mitigation (ROEM)  (?) to mitigate errors arising from qubit measurement, and dynamical decoupling (DD)  (?, ?, ?, ?) to mitigate errors arising from quantum gates. We employed the implementation of ROEM and DD available on the Runtime library of Qiskit  (?), through the Sampler primitive. DD is implemented by sequences of $X$ control pulses, whose effect is to protect qubits from decoherence due to low-frequency system-environment coupling. Here, we applied sequences of two $X$ pulses (as in Ramsey echo experiments) to idle qubits.
@@
Table S1: Details of hardware simulations. For each active space (column 1 from the left), we list the details of the LUCJ quantum circuit executed on hardware, specifically: its number of qubits (q) both for the Jordan-Wigner encoding (JW) and the total number of qubits used including auxiliary ones (Tot.) (column 2) and count of quantum operations (circuit depth $d$, number of $\mathsf{CNOT}$ and single-qubit $\mathsf{u}$ gates, column 3), the device used (column 4) and the qubit layout chosen (column 5, where $[a\rightarrow b]$ indicates the shortest path between qubits $a$ and $b$ in the device topology, also visible in Fig. S8). In columns 6 and 7 we report a 95% confidence interval for the fraction $p_{N}^{\mathrm{hw}}$ of configurations with the correct particle number and the corresponding value for configurations with uniform probability distribution (see the text for the definition of these quantities). In column 8, we list the number of noisy configurations sampled from quantum hardware.
Figure S8: Left to right: schematics of the processors used to carry out experiments for the N2 molecule with 6-31G basis, the N2 molecule with cc-pVDZ basis, the [2Fe-2S] cluster, and the [4Fe-4S] cluster, using local unitary cluster Jastrow (LUCJ) quantum circuits. Qubits used in the calculation are shown in red (for qubits associated with $\alpha$ spin-orbitals), blue for qubits associated with $\beta$ spin-orbitals), and green (for auxiliary qubits).
3.1.2 Classical resource details
The SCI-based eigenstate solver used by SQD is the one implemented in PySCF  (?, ?) for the N2 (6-31G, cc-pVDZ), and [2Fe-2S] systems. For the [4Fe-4S] molecule we use the SCI-based eigenstate solver implemented in DICE  (?, ?) because of its ability to run across multiple computer nodes.
@@ 3.1.2 Classical resource details
4 Additional experimental results
@@
This section contains the results from the experiments that are run on the quantum processors, as described in the previous section. We show a collection of investigations that are complementary to those shown in the main text.
@@
We first visualize the prediction individual wave function components obtained by SQD with configuration recovery in a system size where we have access to the exact ground state wave function amplitudes (N2 in the 6-31G basis).
@@
4.1 Visualization of the wave function amplitudes
@@
Panel (b) in Fig. S9 shows a comparison between the exact ground state wavefunction amplitudes and the wavefunction amplitudes obtained by the estimator, for different points in the dissociation curve. The agreement is exceptional for the larger wavefunction amplitudes and slowly deteriorates in the description of the tails of the wavefunction. We also observe that for larger bond lengths, where higher levels of static correlation are present, the agreement in the tails is worse than for bond lengths closer to the equilibrium geometry. This is to be expected for two reasons. First, for larger bond lengths the wavefunction amplitudes are less concentrated than close to equilibrium. As discussed in Sec. 1.1, the accuracy of the estimator using a SCI-based eigenstate solver deteriorates as the wavefunction becomes less concentrated. The second reason is that for bond lengths close to equilibrium, the wavefunction has a strong mean field character, but not for larger bond lengths. This directly impacts the shape of the reference spin-orbital occupancy ${\bf{n}}$ that is used to recover configurations. When the wavefunction is mean-field dominated, ${\bf{n}}$ takes the shape of a sharp step function shape with all spin-orbital average occupancies close to 0 or 1. One possible effect of static correlations is to soften the step function, moving average occupancies of some spin-orbitals away from 0 or 1. The recovery of configurations used in this work is more effective when the components of ${\bf{n}}$ are close to 0 or 1.
@@
In summary, SQD with configuration recovery produces accurate representations of the ground state wavefunction (on experimental data), even at an individual amplitude level. Less concentrated wavefunctions tend to challenge the procedure more, both because the accuracy of the SCI solver is decreased, and because the configuration recovery becomes less effective due to softer profiles of ${\bf{n}}$.
@@
Figure S9: Visualization of the wavefunctions produced by SQD in the dissociation of N2 (6-31G), and the comparison against the exact wavefunctions. (a) Shows the energy and energy error as a function of the bond length for the estimator with configuration recovery run from measurement outcomes from the quantum processor. Classical approximate methods are shown for reference. (b) Shows the wavefunction amplitudes $|c_{{\bf{x}}}|^{2}$ for all possible electronic configurations ${\bf{x}}$. The configurations in the horizontal axis are ordered according to the magnitude of the amplitudes in the exact wavefunction. The blue solid lines correspond to the exact ground state amplitudes while the orange dots correspond to the wavefunction amplitudes obtained from the estimator with configuration recovery. Different points in the dissociation curve are shown and labelled A, B, C, D, following the same encoding as in panel (a).
4.2 Scaling and runtime
@@
We begin this section by studying the improvement of the accuracy of SQD with the number $d$ of determinants used for the subspace expansion and diagonalization. It is clear that larger $d$ yields more expressivity in the variational wavefunction, since we are increasing the possible size of the support of the state. It is also clear that increasing $d$ increases the HPC quantum estimator runtime since both the projection and diagonalization processes become more computationally intensive. To show the improvement of the accuracy with $d$ we consider the dissociation of N2 (cc-pVDZ), where the measurement outcomes are generated by the quantum processor. The estimator is applied to $10^{5}$ measurement outcomes and we take $K=10$ batches of samples and report $\min_{k}\left(E^{(k)}\right)$ for the potential energy surface for different values of $d$. Ten self-consistent iterations are considered for the recovery of configurations.
@@
Fig. S10 shows the potential energy surfaces obtained by the estimator with different values of $d$, the number of determinants used for the subspace expansion and diagonalization. Increasing the value of $d$ improves the results both qualitatively and quantitatively. For the smaller values of $d$, the potential energy surface shows unphysical oscillations when the bond length is large. For larger values of $d$, not only the energy values are lower, but also the oscillations in the energy as a function of the bond length are significantly decreased in amplitude. Improvement of the ground state approximation with an increasing amount of computational resources is also observed in the energy-variance analyses of the main text (Fig. 4), and the supplementary materials (Figs. S19 and S20).
@@
Figure S10: Improvement of the estimator accuracy (with configuration recovery) by increasing the number of determinants for the subspace expansion and diagonalization. Potential energy surface of the N2 molecule (cc-pVDZ) obtained from the HPC quantum estimator estimator using measurements outcomes from the quantum processor. The energies reported correspond to the lowest energy amongst the K batches of configurations, i.e. ${\min_{k}\left(E^{(k)}\right)}$. Different subspace dimensions $d$ are considered, as indicated in the legend. Different classical methods are shown for reference.
@@
To conclude this section, we study the dependence of the runtime and accuracy of the estimator in the [4Fe-4S] system as a function of the amount of classical compute. The classical compute can be increased by increasing the size of the subspace dimension $d$, or by collecting more batches of configurations $K$. The DICE library allows to distribute the subspace diagonalization across a number of independent computer nodes. It is expected that more classical nodes will yield faster runtimes per diagonalization. For this experiment the Heron processor Montecarlo is used, with $\left|\mathcal{X}\right|=3163742$ collected measurement outcomes, and $p_{N}^{hw}=5.4\cdot 10^{-4}$.
@@
Figure S12: Accuracy and runtime of SQD for the [4Fe-4S] iron-sulfur cluster. (a) Runtime of a single subspace projection and diagonalization as a function of the number of nodes that are used to distribute the calculation. Different colors correspond to different subspace dimensions. The smallest number of nodes per curve corresponds to the limit where the calculation runs out of memory on a single node. The subspace projection and diagonalization are handled by the DICE library. The horizontal line shows a fixed runtime of 1.44 hours for the different subspace sizes. (b) Statistical convergence of SQD as the number of batches of samples $K$ is increased. The vertical axis shows the the average minimum energy obtained by a fixed number of batches of samples. Each batch of samples is run on fixed number of nodes and the diagonalization for the different batches are run in parallel. The number of nodes for the diagonalization of a single batch is chosen to obtain a fixed runtime of 1.44 hours for all of the different subspace dimensions $d$. The horizontal axis shows the total number of nodes corresponding to different numbers of batches. The largest number of batches corresponds to $K=100$.
4.3 Effect of orbital optimizations in the dissociation of N2 (cc-pVDZ)
We apply orbital optimizations (OO) to improve the accuracy of SQD, in the study of the dissociation of N2 (cc-pVDZ). The estimator with configuration recovery and its orbital-optimized counterpart are applied to the same set of noisy measurement outcomes $\tilde{\mathcal{X}}$ obtained from the quantum processor. We use $K=10$ batches of samples in each case and $N_{\textrm{SCF}}=10$ iterations of the alternation between the optimization of $\kappa$ and running of the eigenstate solver to update $c_{\bf{x}}^{(k)}$. After each $c_{\bf{x}}^{(k)}$ update, $5000$ iterations of gradient descent with momentum are used to optimize $\kappa$. Before applying orbital optimizations, the configuration recovery procedure is run for 10 iterations on the reference basis (molecular orbitals). Various sizes $d$ of the batches are considered.